# Grounding Natural Language to SQL Translation with Data-Based Self-Explanations

### v1

@@ I Introduction
Challenges. While significant advancements in enhancing overall accuracy, current end-to-end models face persistent challenges in producing desired quality output during their initial attempt, owing to the treatment of language translation as a “one-time deal”. Figure 1 shows the translation accuracy (defined as query execution result equivalence) on the Spider [7] benchmark, with varied beam outputs for two Seq2seq-based models [^2] Most existing Seq2seq-based NL2SQL translation models utilize beam search decoding method to maintain a list of top-k best outcomes. (i.e., Picard [16], and Resdsql [17]) or diverse chat completion choices for two LLM-based models [^3] LLMs often use a specific parameter to generate various responses. Refer to https://platform.openai.com/docs/api-reference/chat/create#chat-create-n. (i.e., Gpt-3.5-turbo and the state-of-the-art Dailsql model [22]). As indicated by the plateauing accuracy (below 80%) observed when the beam size or the number of chat completions is set to $1$, they may fail to generate best-quality translations in their initial attempts. However, expanding the search space through wider beams or more chat completions steadily improves accuracy, without necessitating modifications to the underlying model architectures. This example shows that end-to-end models may benefit from broader exploration options to enhance translation quality over successive attempts.
@@
Figure 2: An example with simplified database (Flight (left) and Aircraft (right) tables) from the Spider benchmark.
@@
[height=1.0em]thinking faceSQL2NL back-translation for Self-Explanations? A natural way to generate self-provided feedback (explanations) in NL2SQL translation may involve its reverse process, namely SQL2NL. This entails establishing an NL-to-SQL-to-NL translation lifecycle, a concept analogous to back-translation [27, 28]. Several studies have explored this technique [29, 30, 31, 32] to refine initial SQL outputs, but we posit that solely using this “simple” back-translation to generate feedback may be limited, as it lacks additional contextual information beyond the NL and SQL components. Consider an intuitive example in Figure 2, which shows an NL query alongside a translation from an existing model. By simply using the SQL2NL technique, however, incorrect “positive” feedback is generated: the meaning of the explanation seems to align with the initial NL query, whereas the underlying SQL query is deemed incorrect (i.e., an incorrect aggregation function is used in the SELECT clause). Thus, such a back-translated explanation may not adequately serve as valid feedback for the NL2SQL process.
@@
Our Methodology. Cyclesql proposes a plug-and-play framework that facilitates seamless integration with existing end-to-end NL2SQL models. By integrating with a conventional NL2SQL process, Cyclesql 1 first rewrites the translated SQL query to retrieve provenance information from the underlying database for a specific query result. 2 Following this, the provenance is enriched by annotating the semantics of the translated SQL query to its different parts. Next, 3 a data-grounded NL explanation is interpreted from the enriched provenance to rationalize the query result by leveraging text generation techniques [33, 34, 35, 26]. Here, given that the NL explanation integrates both data-level (provenance) and operation-level (query) semantics, they possess significant potential to function as self-provided feedback for various underlying NL2SQL translation models. Finally, 4 the generated NL explanation is used to validate the correctness of the underlying translation iteratively, ultimately yielding a more reliable translation outcome.
@@
To assess the effectiveness of Cyclesql, we conduct a comprehensive experimental evaluation on the widely-used benchmark Spider, and its three variants, namely Spider-realistic [36], Spider-syn [37], and Spider-dk [38], as well as ScienceBenchmark [30], by applying Cyclesql to six contemporary end-to-end NL2SQL models. The results show that Cyclesql can consistently enhance the translation accuracy of all the models. Notably, by applying Cyclesql to Resdsql (with T5-3b scale), Cyclesql obtains a translation accuracy of 82.0% on the validation set and 81.6% on the test set of Spider benchmark, achieving best-reported result among the leading Seq2seq-based models on Spider leaderboard [^4] https://yale-lily.github.io/spider. Moreover, a qualitative evaluation, including a case study and user study, is conducted to show that the explanations generated by Cyclesql can also greatly improve user interpretability in the black-box NL2SQL process.
@@
    Rich Explanations for NL2SQL. The NL explanations generated by Cyclesql, incorporating not only the semantics from the query surface but also the semantics from the data instance, provide insightful information for users to understand the black-box NL2SQL translation process.
@@ II-A NL2SQL models
Seq2seq-based NL2SQL Models. A Seq2seq-based NL2SQL model typically adheres to the encoder-decoder learning framework [39] to translate NL queries to the corresponding SQL queries. Given an input NL query $X=\{x_{1},x_{2},\cdots,x_{n}\}$ and a database schema $S=\left\langle T,C\right\rangle$ that consists of a set of tables $T$ and a set of corresponding columns $C$, the model uses an encoder to compute a contextual representation c by jointly embedding the NL query $X$ with schema $S$. Afterward, an auto-regressive decoder is used to compute a distribution $P(Y\mid\textbf{c})$ over the SQL programs $Y=(y_{1},\cdots,y_{\mid Y\mid})$.
LLMs as NL2SQL Models. In light of the recent advancements in LLMs, current research [40, 41] endeavors to employ LLMs as NL2SQL models without fine-tuning. Given an NL query $X$ and a prompt $P$ as input, an LLM can be utilized to generate the corresponding SQL query $Y$. Depending on the prompting technique utilized, such as zero-shot, few-shot prompting, or in-context learning, the prompt $P$ can include text instructions [41], translation demonstrations [20, 22] or reasoning chains [19, 42].
@@ II-B Provenance
Provenance elucidates the origin and history of data throughout its lifecycle. Research has studied provenance semantics for various query classes in databases [43, 44]. The most common forms of database provenance describe relationships between data in the source and its corresponding output [45]. This includes explaining where output data originated in the input (commonly known as where-provenance) [46, 47], revealing inputs that explain why an output record was produced (referred to as why-provenance) [48, 47], or providing a detailed account of how an output record was generated (referred to as how-provenance) [49]. This work adopts why-provenance, which functions as evidence within a query, comprising a subset of tuples that guarantees the presence of a specific tuple in the output.
@@ III-A Natural Language Explanation
Explanations for database query answers can differ across various dimensions, encompassing clarifying unexpected query outcomes [50], explaining the rationale behind generating intriguing results from a query [51], and justifying how an answer aligns with the query’s criteria [33, 34]. Each of these dimensions serves a distinct purpose in query explanation.
@@
Example 1. The following shows an NL explanation generated by Cyclesql (details of the generation process will be explained in late sections), which mostly includes two intermediate reasoning steps specifying the details of the reasoning process from the initial NL query to the query result based on the translated SQL query in the example shown in Figure 2.
@@ III-B Feedback Loop in Cyclesql
A high-level view of Cyclesql is presented in Figure 3. When seamlessly integrated into the end-to-end NL2SQL process of a designated translation model, the fundamental operational sequence of Cyclesql unfolds as follows:
@@
Data Tracking. In the initial stage (illustrated in Figure 3- 1), Cyclesql entails tracing the lineage of the query result obtained from the execution of a SQL query sampled from an underlying translation model, intending to gather provenance information. In Cyclesql, the tracing process is implemented by rewriting the sampled SQL query to retrieve all relevant data-level details from the underlying database.
For example, regarding the query result shown in Figure 2, its provenance corresponds to tuples $F2$, $F3$, where the column “aid” is equal to $3$ in “Flight” table, as below,
@@
Semantics Enrichment. In the following stage (depicted in Figure 3- 2), Cyclesql further enriches the provenance information by integrating the operation-level semantics from the translated SQL query into its different parts. Since the translated SQL query captures the semantics from the initial NL query, this integration enables the reflection of user query semantics across different data segments of provenance. Cyclesql achieves this by decomposing the SQL query into a set of query units and associating the semantics of each query unit with relevant parts of data provenance.
Continuing the example in Figure 2, the column “name” in the provenance table is enriched with the filtering semantics from the WHERE operation, and the whole table is associated with the count semantics from SELECT to attempt to reflect the “all flight numbers” semantics in the original NL query.
Explanation Generation. In the subsequent stage (depicted in Figure 3- 3), Cyclesql interprets the enriched provenance information into an NL expression, generating a detailed explanation for the given query result. Inspired by the success of text generation techniques, Cyclesql utilizes a simple but effective rule-based method to synthesize the NL explanation mechanically from the given enriched provenance information.
For instance, the NL explanation produced by Cyclesql for query result in Figure 2 can be seen in Example $1$ in Section III-A, which interprets the meaning of number $2$.
Translation Verification. In the final stage (shown in Figure 3- 4), Cyclesql employs the generated NL explanation of the given query result as self-provided feedback to verify the accuracy of the underlying NL2SQL translation. To accomplish this, Cyclesql formulates the translation validation process as a natural language inference (NLI) task in the NLP field and uses a deep-learning model to determine whether the NL explanation entails the original NL query, thereby validating the correctness of the translation. Cyclesql iterates through the above stages until validation succeeds, ultimately leading to an improved translation outcome.
Continuing the example in Figure 2, Cyclesql may recognize that expressions such as “there are 2 flights in total” in the explanation are not aligned with the “all flight numbers” semantics of the initial NL query, thereby the translated SQL query is validated as incorrect. In response, Cyclesql solicits another sampled SQL query from the underlying translation model, repeating this process until the validation succeeds.
@@ IV-A Provenance via Query Rewriting
  • Rule 3 (Aggregation Deconstructive Rule). Since aggregation functions (and the GROUP BY clause) may collapse input data rows, they inherently conceal the ability to track the provenance. Therefore, we invert the potential transformations by simply eliminating aggregation functions (and GROUP BY clause) to unravel the lineage of the data [^5] Note that these operations are temporarily excluded for query rewriting purpose. They will be revisited in the following phase for query semantics annotation. Further details can be found in the late section..
@@
Example 2. Figure 4 illustrates the query rewriting process applied to the translated SQL query presented in Figure 2. The rewritten SQL query is transformed by applying Rule $2$ to retrieve the column from the condition and Rule $3$ to eliminate the aggregation by replacing it with relevant columns.
It is important to note that as Cyclesql attempts to trace the provenance information of a given query result by retrieving the relevant data records, the absence of a result from an SQL query may indicate a lack of corresponding provenance information. Therefore, Cyclesql skips this step for empty-result queries and directly uses the operation-level semantics (refer to the subsequent section) to generate the explanation instead. Exploring more effective strategies for managing empty-result queries within the context of Cyclesql could pose an interesting research problem.
@@ IV-B Semantics Enrichment via Annotating
The main observation arising from the provenance enrichment process is that while provenance information can reveal details about the query result derived from the data, it is insufficient for providing insights into why the query result satisfies the query criteria [51]. To give a flavor of the explanation that may be produced by only leveraging the provenance information, an NL explanation is exemplified as follows for the example in Figure 4,
@@
Although the explanation interprets the details about those flights with “Airbus A340-300” aircraft, it does not clarify how result number $2$ is derived. This lack of explicit information may impede Cyclesql from getting a valid explanation suitable for validating the underlying translation afterward.
@@
Example 3. Figure 5 illustrates the enriched provenance table retrieved in Figure 4, where the table is enriched by one annotation (i.e., 2) from the original translated query, along with one annotation (i.e., 1) derived from the WHERE clause.
@@ IV-C Explanation Generation via Graphing
More specifically, Cyclesql initiates the process by generating a brief text summary of the query, outlining details such as the number of columns and rows in the returned result set. Next, Cyclesql partitions the enriched provenance table of a specific query result into distinct segments, proceeding with the incremental construction of a provenance graph. A provenance graph, denoted as $G_{p}(V_{p},E_{p})$, is a directed graph where nodes in $V_{p}$ represent provenance elements, such as the table element “Flights”, the column element “filghtNo” and the value element “1377”. Each edge in $E_{p}$ is associated with a type, indicating a specific relationship between two elements. For example, as for the provenance table in Figure 5, $G_{p}$ has a joint table node “Airlines-Flights”, which has a “hasAttribute” type edge that connects with the column node “uid”; the column node “uid”, in turns, has a “hasValue” type edge that connects with the value node “12”. Then, Cyclesql assigns the query annotations we obtained in Section IV-B as the semantics labels to each corresponding element. Finally, Cyclesql traverses the graph $G_{p}$ to generate NL phrases for each provenance element with its associated labels. The NL phrases (and the initially generated summary) are then concatenated to generate the NL explanation by using some descriptive expressions (e.g., ‘so’, ‘for’, etc.). We present the explanation generation process in Algorithm 1.
@@
Join-related Semantics. We observe that the semantics expressed in provenance may become much more abstract when derived from join-related operations [35]. For example, the provenance table shown in Figure 5 represents a new “table” derived from the join operation of the underlying SQL query, and the semantics of the aggregation-related semantics (i.e., count semantics) become abstract after the table joining.
@@
When a join relation appears in a query, we convert it into the graph structure as described above and then check for isomorphism with any pre-defined graphs to determine the corresponding join semantics. If no isomorphism is found, we use the associated table names to represent the join semantics instead. Figure 6 shows an example of a three-table query, where the “Singer_in_concert” table has two foreign keys linked to the “Concert” and “Singer” tables. When matched with pre-defined graphs, an isomorphism graph (i.e., the three-node graph with subject-relationship-object semantics) is identified, enabling the join semantics to be instantiated as “singer with concert” using the specific table names.

### v2

@@ I Introduction
Challenges. While significant advancements in enhancing overall accuracy, current end-to-end models face persistent challenges in producing desired quality output during their initial attempt, owing to the treatment of language translation as a “one-time deal”. Fig. 1 shows the translation accuracy (defined as query execution result equivalence) on the Spider [7] benchmark, with varied beam outputs for two Seq2seq-based models [^2] Most existing Seq2seq-based NL2SQL translation models utilize beam search decoding method to maintain a list of top-K best outcomes. (i.e., Picard [16], and Resdsql [17]) or diverse chat completion choices for two LLM-based models [^3] LLMs often use a specific parameter to generate various responses. Refer to https://platform.openai.com/docs/api-reference/chat/create#chat-create-n. (i.e., Gpt-3.5-turbo and the state-of-the-art Dailsql model [22]). As indicated by the plateauing accuracy (below 80%) observed when the beam size or the number of chat completions is set to $1$, they may fail to generate best-quality translations in their initial attempts. However, expanding the search space through wider beams or more chat completions steadily improves accuracy, without necessitating modifications to the underlying model architectures. This example shows that end-to-end models may benefit from broader exploration options to enhance translation quality over successive attempts.
@@
Figure 2: A simplified example from the Spider benchmark.
@@
[height=1.0em]thinking faceSQL2NL back-translation for Self-Explanations? A natural way to generate self-provided feedback (explanations) in NL2SQL translation may involve its reverse process, namely SQL2NL. This entails establishing an NL-to-SQL-to-NL translation lifecycle, a concept analogous to back-translation [27, 28]. Several studies have explored this technique [29, 30, 31, 32] to refine initial SQL outputs, but we posit that solely using this “simple” back-translation to generate feedback may be limited, as it lacks additional contextual information beyond the NL and SQL components. Consider an intuitive example in Fig. 2, which shows an NL query alongside a translation from an existing model. By simply using the SQL2NL technique, however, incorrect “positive” feedback is generated: the meaning of the explanation seems to align with the initial NL query, whereas the underlying SQL query is deemed incorrect (i.e., an incorrect aggregation function is used in the SELECT clause). Thus, such a back-translated explanation may not adequately serve as valid feedback for the NL2SQL process.
@@
Our Methodology. We propose a plug-and-play framework that facilitates seamless integration with existing end-to-end NL2SQL models. By integrating with a conventional NL2SQL process, Cyclesql 1 first rewrites the translated SQL query to retrieve provenance information from the underlying database for a specific query result. 2 Following this, the provenance is enriched by annotating the semantics of the translated SQL query to its different parts. Next, 3 a data-grounded NL explanation is interpreted from the enriched provenance to rationalize the query result by leveraging text generation techniques [33, 34, 35, 26]. Here, given that the NL explanation integrates both data-level (provenance) and operation-level (query) semantics, they possess significant potential to function as self-provided feedback for various NL2SQL models. Finally, 4 the generated explanation is used to validate the correctness of the underlying translation iteratively, ultimately yielding a more reliable translation outcome.
@@
To assess the effectiveness of Cyclesql, we conduct a comprehensive experimental evaluation on the widely-used benchmark Spider, and its three variants, namely Spider-realistic [36], Spider-syn [37], and Spider-dk [38], as well as ScienceBenchmark [30], by applying Cyclesql to seven contemporary end-to-end models. The results show that Cyclesql can consistently enhance the translation accuracy of all the models. Notably, by applying Cyclesql to Resdsql (with T5-3b scale), Cyclesql obtains a translation accuracy of 82.0% on the validation set and 81.6% on the test set of Spider benchmark, achieving best-reported result among the leading Seq2seq-based models on Spider leaderboard [^4] https://yale-lily.github.io/spider. Moreover, a qualitative evaluation, including a case study and user study, is conducted to show that the explanations generated by Cyclesql can also greatly improve user interpretability in the black-box NL2SQL process.
@@
    Rich Explanations for NL2SQL. The NL explanations generated by Cyclesql, which incorporate not only the semantics of the query surface but also the semantics from the data instance, provide insightful information for users to understand the black-box NL2SQL translation process.
@@ II-A NL2SQL models
Seq2seq-based NL2SQL Models. A Seq2seq-based NL2SQL model typically adheres to the encoder-decoder learning framework [39] to translate NL queries to the corresponding SQL queries. Given an input NL query $X=\{x_{1},x_{2},\cdots,x_{n}\}$ and a database schema $S=\left\langle T,C\right\rangle$ that consists of a set of tables $T$ and a set of corresponding columns $C$, the model uses an encoder to compute a contextual representation c by jointly embedding the NL query $X$ with schema $S$. Afterward, an autoregressive decoder is used to calculate a distribution $P(Y\mid\textbf{c})$ over the SQL programs $Y=(y_{1},\cdots,y_{\mid Y\mid})$.
LLMs as NL2SQL Models. In light of recent advancements in LLMs, current research [40, 41] endeavors to employ LLMs as NL2SQL models without fine-tuning. Given an NL query $X$ and a prompt $P$ as input, an LLM can be utilized to generate the SQL query $Y$. Depending on the prompting technique utilized, such as zero-shot, few-shot prompting, or in-context learning, the prompt $P$ may include instructions [41], demonstrations [20, 22] or reasoning chains [19, 42].
@@ II-B Provenance
Provenance elucidates the origin and history of the data throughout its lifecycle. Research has studied provenance semantics for various query classes in databases [43, 44]. The most common forms of database provenance describe the relationships between the data in the source and its corresponding output [45]. This includes explaining where the output data originated in the input (commonly known as where-provenance) [46], revealing inputs that explain why an output record was produced (referred to as why-provenance) [47, 48], or providing a detailed account of how an output record was generated (referred to as how-provenance) [49]. This work adopts why-provenance, which functions as evidence within a query, comprising a subset of tuples that guarantees the presence of a specific tuple in the output.
@@ III-A Natural Language Explanation
Explanations of database query answers can differ across various dimensions, encompassing clarifying unexpected query outcomes [50], explaining the rationale behind generating intriguing results from a query [51], and justifying how an answer aligns with the query criteria [33, 34]. Each of these dimensions serves a distinct purpose in query explanation.
@@
Example 1. The following shows an NL explanation generated by Cyclesql (details of the generation process will be explained in late sections), which mostly includes two intermediate reasoning steps specifying the details of the reasoning process from the initial NL query to the query result based on the translated SQL query in the example shown in Fig. 2.
@@ III-B Feedback Loop in Cyclesql
A high-level view of Cyclesql is presented in Fig. 3. When seamlessly integrated into the end-to-end NL2SQL process of a designated translation model, the fundamental operational sequence of Cyclesql unfolds as follows:
@@
Data Tracking. In the initial stage (illustrated in Fig. 3- 1), Cyclesql entails tracing the lineage of the query result obtained from the execution of a SQL query sampled from an underlying translation model, intending to gather provenance information. In Cyclesql, the tracing process is implemented by rewriting the sampled SQL query to retrieve all relevant data-level details from the underlying database.
For example, regarding the query result shown in Fig. 2, its provenance corresponds to tuples $F2$, $F3$, where the column “aid” is equal to $3$ in “Flight” table, as below,
@@
Semantics Enrichment. In the following stage (depicted in Fig. 3- 2), Cyclesql further enriches the provenance information by integrating the operation-level semantics from the translated SQL query into its different parts. Since the translated SQL query captures the semantics from the initial NL query, this integration enables the reflection of user query semantics across different data segments of provenance. Cyclesql achieves this by decomposing the SQL query into a set of query units and associating the semantics of each query unit with relevant parts of data provenance.
Continuing the example in Fig. 2, the column “name” in the provenance table is enriched with the filtering semantics from the WHERE operation, and the whole table is associated with the count semantics from SELECT to attempt to reflect the “all flight numbers” semantics in the original NL query.
Explanation Generation. In the subsequent stage (depicted in Fig. 3- 3), Cyclesql interprets the enriched provenance information into an NL expression, generating a detailed explanation for the given query result. Inspired by the success of text generation techniques, Cyclesql utilizes a simple but effective rule-based method to synthesize the NL explanation mechanically from the given enriched provenance information.
For instance, the NL explanation produced by Cyclesql for query result in Fig. 2 can be seen in Example $1$ in Section III-A, which interprets the meaning of number $2$.
Translation Verification. In the final stage (shown in Fig. 3- 4), Cyclesql employs the generated explanation as self-provided feedback to verify the accuracy of the underlying NL2SQL translation. To accomplish this, Cyclesql formulates the validation process as a natural language inference (NLI) task and uses a deep-learning model to determine whether the NL explanation entails the original NL query, thereby validating the correctness of the translation. Cyclesql iterates through the above stages until validation succeeds, ultimately leading to an improved translation outcome.
Continuing the example in Fig. 2, Cyclesql may recognize that expressions such as “there are 2 flights in total” in the explanation are not aligned with the “all flight numbers” semantics of the initial NL query, thereby the translated SQL query is validated as incorrect. In response, Cyclesql solicits another sampled SQL query from the underlying translation model, repeating this process until the validation succeeds.
@@ IV-A Provenance via Query Rewriting
  • Rule 3 (Aggregation Deconstructive Rule). Since aggregation functions (and the GROUP BY clause) may collapse input data rows, they inherently conceal the ability to track the provenance. Therefore, we invert the potential transformations by simply eliminating aggregation functions (and GROUP BY clause) to unravel the lineage of the data [^5] Note that these operations are temporarily excluded for query rewriting purpose. They will be revisited in the following phase for query semantic annotation. More details can be found in the later section..
@@
Example 2. Fig. 4 illustrates the query rewriting process applied to the translated SQL query presented in Fig. 2. The rewritten SQL query is transformed by applying Rule $2$ to retrieve the column from the condition and Rule $3$ to eliminate the aggregation by replacing it with relevant columns.
It is important to note that as Cyclesql attempts to trace the provenance information of a given query result by retrieving the relevant data records, the absence of a result from an SQL query may indicate a lack of corresponding provenance information. Therefore, Cyclesql skips this step for empty-result queries and uses the operation-level semantics directly (see the subsequent section) to generate the explanation instead. Exploring more effective strategies for managing empty-result queries within the context of Cyclesql could pose an interesting research problem.
@@ IV-B Semantics Enrichment via Annotating
The main observation arising from the provenance enrichment process is that while provenance can reveal details about the query result derived from the data, it is insufficient to provide insights into why the query result satisfies the query criteria [51]. To give a taste of the explanation that can be produced using provenance information alone, an explanation is exemplified as follows for the example in Fig. 4,
@@
Although the explanation interprets the details about those flights with “Airbus A340-300” aircraft, it does not clarify how the resulting number $2$ is derived. This lack of explicit information may impede Cyclesql from getting a valid explanation suitable for validating the translation afterward.
@@
Example 3. Fig. 5 illustrates the enriched provenance table retrieved in Fig. 4, where the table is enriched by one annotation (i.e., 2) from the original translated query, along with one annotation (i.e., 1) derived from the WHERE clause.
@@ IV-C Explanation Generation via Graphing
More specifically, Cyclesql initiates the process by generating a brief text summary of the query, outlining details such as the number of columns and rows in the returned result set. Next, Cyclesql partitions the enriched provenance table of a specific query result into distinct segments, proceeding with the incremental construction of a provenance graph. A provenance graph, denoted as $G_{p}(V_{p},E_{p})$, is a directed graph where nodes in $V_{p}$ represent provenance elements, such as the table element “Flights”, the column element “filghtNo” and the value element “1377”. Each edge in $E_{p}$ is associated with a type, indicating a specific relationship between two elements. For example, as for the provenance table in Fig. 5, $G_{p}$ has a joint table node “Airlines-Flights”, which has a “hasAttribute” type edge that connects with the column node “uid”; the column node “uid”, in turns, has a “hasValue” type edge that connects with the value node “12”. Then, Cyclesql assigns the query annotations we obtained in Section IV-B as the semantics labels to each corresponding element. Finally, Cyclesql traverses the graph $G_{p}$ to generate NL phrases for each provenance element with its associated labels. The NL phrases (and the initially generated summary) are then concatenated to generate the NL explanation by using some descriptive expressions (e.g., ‘so’, ‘for’, etc.). We present the explanation generation process in Algorithm 1.
@@
Join-related Semantics. We observe that the semantics expressed in provenance may become much more abstract when derived from join-related operations [35]. For example, the provenance table shown in Fig. 5 represents a new “table” derived from the join operation of the underlying SQL query, and the semantics of the aggregation-related semantics (i.e., count semantics) become abstract after the table joining.
@@
When a join relation appears in a query, we convert it into the graph structure as described above and then check for isomorphism with any pre-defined graphs to determine the corresponding join semantics. If no isomorphism is found, we use the associated table names to represent the join semantics instead. Fig. 6 shows an example of a three-table query, where the “Singer_in_concert” table has two foreign keys linked to the “Concert” and “Singer” tables. When matched with pre-defined graphs, an isomorphism graph (i.e., the three-node graph with subject-relationship-object semantics) is identified, enabling the join semantics to be instantiated as “singer with concert” using the specific table names.